# Military AI Needs Technically-Informed Regulation to Safeguard AI Research and its Applications

### v1

@@ 1 Introduction
Artificial intelligence (AI), also called machine learning (ML), is already embedded in a growing number of deployed weapon systems. These include uncrewed aerial vehicles, uncrewed surface vessels and submarines, and battlefield coordination platforms that assist with real-time targeting and decision support [1, 2, 3, 4, 5, 6, 7]. The rapid deployment of these AI-augmented systems has outpaced existing regulatory frameworks. Governance models for conventional lethal autonomous weapon systems (LAWS)—typically rule-based and non-adaptive—were not designed to address challenges posed by modern AI [8, 9, 10, 11]. These challenges include opacity (hard-to-interpret decision processes), adaptivity (behavioral shifts in response to new inputs), and post-deployment drift (performance degradation over time) [2, 12, 13, 14].
Yet in both policy and ML communities, distinctions between conventional LAWS and modern AI-powered systems remain poorly defined—and are absent from existing treaties, principles, and deployment norms. These systems may misclassify targets under unfamiliar conditions, escalate conflicts due to unpredictable outputs, or obscure human accountability through black-box decisions—all while using AI models originally developed in civilian contexts (see Table 1).
As these systems become more central to warfare, civilian AI advances gain national security relevance, and governments may restrict the movement of people, research, and AI services to gain military advantage [12, 15]. Given the risks that reckless development of AI weapons pose, we take the position that modern AI-powered weapon systems—what we call AI-LAWS—require distinct and novel regulations to avoid harms to AI research and the AI industry. Effective regulation must consider the underlying ML technologies, and AI experts from diverse fields must be directly involved in the regulatory process.
To motivate this, we first review the current state of military AI development and policy in section 2, highlighting the range of systems publicly announced and already deployed. Next, we introduce a functional definition of AI-LAWS based on AI/ML involvement in targeting, advising, or decision-making (section 3), which we use to identify systems that pose novel risks and require distinct regulatory consideration. Lastly, we propose several technically-informed policy directions grounded in this definition to address the risks AI-LAWS pose to both AI research and global security (section 4).
@@ 2 Background
Before outlining our contributions, we provide context by surveying deployed and in-development systems that meet our criteria for AI-LAWS (section 3), with examples listed in Table 2. We then examine the unique risks these systems pose in subsection 2.2, and explain why AI expert involvement is essential. These risks are not theoretical—they are structural features of how AI-LAWS behave in practice, already visible in deployed systems. Performance may degrade in new environments, behaviors may escalate tensions, and lethal decisions may be made by models no one fully understands. We summarize these risks in Table 1. Finally, we review existing regulatory efforts and alternate positions in subsection 2.3, and highlight why they fail to address risks specific to AI-LAWS.
@@ 2.1 Existing AI-LAWS
AI-LAWS are already deployed worldwide. Systems using AI to make real-time targeting or coordination decisions have already been demoed and deployed across air, land, sea, and command domains. Table 2 highlights examples with documented AI capabilities, from drone swarms to battlefield planning platforms. Many of these systems rely on models and tools developed in civilian research settings. Yet, AI researchers often remain unaware that their work may shape combat outcomes—or be deployed without rigorous validation or constraint [15, 16]. Some AI-LAWS have already seen combat [15, 10]. The Russian Lancet loitering munition, for example—a drone that independently identifies and crashes into targets—reportedly includes AI-based autonomous targeting capabilities and has seen wide deployment in Ukraine [17, 18, 19]. AI-LAWS span a wide range of platforms, from disposable drones to naval vessels and advanced aircraft [20] (Table 2). These systems are being developed by both major and minor powers—with many systems being actively deployed [21, 22, 15]. For example, the Estonian Milrem THeMIS UGV, which includes autonomous navigation and potentially targeting, has seen battlefield use in Ukraine [23, 24]. Given the diversity of actors, platforms, and missions, even well-intentioned policies will fall short unless they account for the varied ways AI models are used across military operations [10, 25].
Table 1: Risks introduced by AI-LAWS that meet our oversight criteria. Oversight must be grounded in technical behavior—not system intent—and informed by expertise from the AI research community.
@@
Undetected failures in real-world use Gaps in validation and overtrust in AI outputs make it harder to detect failures during deployment—e.g., in stressful or unfamiliar contexts. Forest-trained targeting model misidentifies desert vehicles [26, 27] and commanders defer to faulty AI planning advice [28, 29].
Opacity and black-box decision-making AI/ML-based targeting systems are difficult for human operators to understand or audit, causing unpredictable behavior in combat. AI-enabled engagement system selects lethal target based on faulty sensors the operator cannot override in time, leading to unintended casualties [13].
Erosion of AI researcher freedom and scientific exchange Classified funding streams and dual-use constraints limit openness, autonomy, and international collaboration. University AI lab begins operating under military classification rules, limiting publication and/or international collaboration [30, 12, 31].
Channeling AI expertise toward military applications Civilian AI researchers are recruited into military programs, often without clear disclosure or opt-out mechanisms. University scientists find their programs absorbed into a defense-funded lab, shifting away from open-source AI research [32, 30].
Acceleration of arms races and regional or global instability Widespread access to AI-LAWS lowers barriers to high-risk escalation. Deploying drone swarms without adequate testing/safeguards, risking conflict escalation [33, 34].
Table 2: Examples of real-world AI-enabled military systems with publicly documented unmanned or AI-centered capabilities. Includes platforms used for targeting, coordination, or decision support across air, land, sea, and command domains. Status reflects 2019–2024 public data: In Dev = In development, Demo = Demonstrated, Deployed = In deployment.
@@
18th Airborne Corps System [35] U.S. Army U.S. - Deployed
Lavender [4] IDF Israel - Deployed
ChatBIT [36] PLA China - Demo
@@ 2.2 AI-LAWS present unique risks
AI-LAWS create distinct risks across military, geopolitical, and institutional domains. AI-LAWS pose novel risks across three overlapping domains—military [10], international relations [11, 2, 37], and institutional (e.g., universities, tech companies) [12, 15, 38, 39]—that differ from those associated with conventional LAWS. First, poorly validated or overtrusted systems can cause battlefield failures—including misidentification, friendly fire, brittle AI-human coordination, and performance gaps between exercises and combat [40, 41, 29, 42, 43]. Second, unpredictable or escalatory behavior can increase the risk of interstate conflict, especially when AI systems are deployed without adequate testing or validation [44, 13, 2, 29]. Third, the rapid pipeline from AI research to military use threatens openness, research direction, and international norms [45, 10, 12, 16]. We summarize these risks in Table 1.
@@
In the military domain, the need to validate weapons performance during peacetime is not new—but the characteristics of modern AI make it more difficult than for conventional systems. Without additional validation, AI-LAWS are more likely to underperform in conflict, often in unintuitive ways. For example, “brittleness” is a well-documented issue in both academic and industrial AI, especially as systems become more advanced and multifunctional. It refers to the often rapid decline in performance when AI is exposed to contexts outside its training data—including variations a human would expect to be irrelevant [46, 47]. For example, an AI-LAWS trained on diverse biomes but validated only in temperate European environments may underperform in tropical or East Asian settings. This risk heightens when iterative tuning focuses on a narrow deployment context, causing unintentional overfitting to the validation setting. As a result, performance suffers in environments the system was trained on but never field-tested in. Such failures can also be difficult to detect— the AI models AI-LAWS rely on are partially opaque, and their decision bases may not be interpretable or auditable with high confidence [8, 3, 14].
@@
For the second domain, beyond battlefield performance, the development and spread of AI-LAWS may destabilize international relations in subtle and unpredictable ways. For example, the unpredictability of modern AI systems can create novel conflict risks if an AI-LAWS is deployed in a diplomatically tense region [13, 2]. An AI-LAWS patrolling a border may protect soldiers, but could misinterpret warning shots as hostile and initiate or recommend return fire. Even without deployment, uncertainty over AI-LAWS battlefield impact has fueled arms-race dynamics, prompting major investments by multiple nations [37, 20]. These investments risk becoming reciprocal responses among rivals—accelerating a competitive arms race. The problem is compounded by technical uncertainty around AI-LAWS capabilities and timelines, and the need to test systems in real world combat [41, 48, 40]. Lacking better insight, policymakers may overestimate their impact to avoid under-investing, while also seeking conflict to test them in. These dynamics raise the risk of escalation even before AI-LAWS are deployed.
@@
Lastly, AI-LAWS pose risks to academic and civilian AI freedoms and norms. Modern AI is distinctive in that the gap between civilian and military applications is unusually small. As a result, civilian advances are easily repurposed for military use, which has national and global security implications. This incentivizes governments to overregulate the process and dissemination of AI research and models, e.g., through publication blackouts and travel restrictions on AI researchers [12]. As civil-military barriers erode, regulatory oversight, export restrictions, and international controls may further limit civilian AI development. Some restrictions—such as controls on AI hardware exports—are already in place, and similar constraints on software and algorithms are likely to follow [49, 50, 51]. These risks are amplified by uncertainty: when policymakers lack clarity about AI’s national security implications, they may over-regulate to hedge against worst-case scenarios. These governance failures are addressed in the next section.
@@ 2.3 Existing policy efforts do not address unique AI-LAWS risks
Despite broad policy attention, current governance frameworks do not account for the technical risks posed by modern AI systems. Nations continue to field military AI, yet no binding standards exist to govern how the systems are designed, validated, or deployed [44]. Most existing LAWS and AI safety frameworks emphasize broad principles—such as human oversight or bias minimization—but lack concrete mechanisms and metrics tailored to how these systems function in practice.
Alongside these awareness issues, the AI research community has been largely isolated from military governance conversations. Initial regulatory efforts were led by defense officials, with limited technical input from the scientific community [52]. Today, oversight of military AI is debated at institutional, national, and international levels [53, 25]—yet those best equipped to evaluate system behavior remain underrepresented. Militaries worldwide have expressed concern over lacking AI expertise to update doctrine and regulation as technology evolves [15, 54, 55, 56]. This disconnect creates both risk and opportunity. To meaningfully address AI-specific failure modes, researchers must be at the table [57]. AI researchers who seek to shape policy must engage with military actors beyond their usual academic spheres [16, 15, 58]. Engagement is essential to ensure that governance reflects real-world system behavior—not just design intentions or policy abstractions.
Regulatory frameworks have not kept pace with failure modes such as opacity, drift, and escalatory misuse [20]. Technical and policy communities alike struggle to establish thresholds for oversight, especially as systems grow more adaptive, autonomous, and complex. Current certification procedures are rooted in conventional systems and offer little clarity on AI-specific metrics or standards [44, 59, 55]. These gaps are compounded by geopolitical concerns: policymakers may resist unilateral regulation for fear of losing strategic advantage in advancing AI technologies [37], and international progress has been limited by misaligned priorities and inconsistent institutional mandates [60].
Efforts to define AI-LAWS—in the United Nations [37], Track II forums [61, 62], or REAIM [63]—have struggled to yield useful standards [64]. Technical experts remain underrepresented in these discussions, and proposed definitions often hinge on extreme thresholds like “full autonomy” or “superhuman learning” [52]. One influential definition, for example, requires that a system expand its capabilities beyond human expectations without any human intervention [37]. While well-intentioned, such thresholds exclude most deployed and near-term systems—and thereby prevent oversight of the systems already in use.
@@ 3 Oversight Criteria for AI-LAWS, Grounded in System Behavior
Effective oversight of AI-LAWS requires moving beyond labels like “autonomous” or “general intelligence” and toward criteria grounded in how systems actually function. Targeting models can misidentify civilian vehicles in unfamiliar terrain [26, 13]. Autonomous patrols may escalate based on ambiguous signals, without human authorization [2, 13]. Command-advising models may drift over time in ways operators cannot detect [3, 43]. These risks are not hypothetical—they mirror failures already observed or anticipated in real-world systems (see Table 1). Oversight must begin with how systems behave, not how they are labeled.
Current frameworks often hinge on vague descriptors—like “human-in-the-loop” or “fully autonomous”—that fail to reflect how systems behave in practice. For example, a loitering munition may shift between human-authorized and AI-executed modes based on mission context [22]. Command-advising systems may shape lethal decisions without directly executing them [9, 3]. Despite posing real governance challenges, such systems risk not qualifying as AI-LAWS under strict definitional thresholds—even when they should.
@@
  • Criterion 1: Use of ML/AI methods (e.g., neural networks) that are necessary for system function, and that pose AI-specific risks such as misidentifying targets, escalating unpredictably, drifting after deployment, or degrading in unfamiliar environments due to limited generalization beyond the training data distribution.
@@
One of the main barriers to regulating conventional LAWS has been disagreement over definitions [65, 64]. According to U.S. submissions to the UN, LAWS—sometimes called “killer robots”—are defined as systems that “can identify, select, and engage targets with lethal force without further intervention by an operator” [66]. Critics note that this definition is overly broad, potentially encompassing naval mines and heat-seeking missiles. The radar-seeking Israeli Harpy loitering munition, deployed in the late 1980s, exemplifies early LAWS already covered by prior regulations as historical models behave predictably with simple tracking functions [65]. Recent advances in AI research have significantly increased the combat power and autonomy of AI-enhanced systems [67, 68, 11, 3, 5, 15]. As AI capabilities were added to LAWS, it became increasingly necessary to distinguish conventional automated weapons from AI-LAWS [22, 25, 59]. While AI has long been used in simple forms for military tasks [69, 33], modern AI-LAWS employ far more advanced ML models, can operate along a continuum of modes from remote-controlled to fully autonomous, and present unique AI-specific risks and failure modes not seen in conventional LAWS.
@@
Oversight of military AI systems has traditionally centered around three narratives. While each raises valid concerns, none is technically grounded enough to address how deployed systems behave—particularly regarding opacity, drift, and brittleness under domain shift. The first, rooted in humanitarian advocacy, calls for bans or strict human-in-the-loop constraints based on the principle of Meaningful Human Control (MHC) [70, 71, 21]. The second emphasizes long-term existential risks from future artificial general intelligence (AGI) [72, 73]. The third—most common in defense policy—treats AI-LAWS as a natural extension of conventional automation and maintains that existing frameworks, such as the U.S. DoD’s Directive 3000.09, are sufficient [59]. This approach, termed Appropriate Human Judgment (AHJ), is embedded in U.S. military doctrine [59, 74]. It emphasizes procedural safeguards—such as system testing, operator training, and robust design—rather than requiring direct human control over every lethal force application. AHJ is often viewed as more pragmatic for increasingly complex systems—such as when operations require faster response times than humans can provide, involve cooperative autonomous swarms, or scale beyond what human oversight can track in real time [7].
@@
Yet all three narratives remain underdefined and difficult to operationalize. MHC lacks measurable criteria for what constitutes “meaningful” control in fast-moving or distributed contexts. AHJ, though more scalable, often fails to define metrics for what constitutes sufficient “judgment”—or to codify how to assess the safety of opaque, adaptive, or continually learning systems [59]. Critics note that AHJ can obscure accountability behind vague assurances of system design—especially when those assurances come from third-party contractors or defense tech firms [13, 2, 16, 43, 75].
@@
Policy limitations have real consequences. Post-deployment drift, overreliance on black-box systems, brittle generalization, and institutional pressure to avoid challenging flawed AI outputs can all occur under either oversight model [20]. Command-advising systems may formally keep humans “in the loop,” yet users often defer to flawed recommendations—especially under time pressure or when authority to intervene is unclear [76, 29]. To be effective, any oversight framework—whether based on MHC, AHJ, or another model—must be paired with rigorous behavioral validation, grounded in AI research expertise. This includes task-specific performance benchmarks, failure audits, stress testing under domain shift, and post-deployment monitoring. These mechanisms do not replace high-level policy principles; they make them testable and enforceable. The U.K.’s 2024 guidance for military AI use offers a promising start, with proposals for AI-specific audits, operator certification, and red-line prohibitions [55]. However, as AI capabilities advance, fixed rules alone cannot keep pace. Governance must evolve alongside systems with flexible frameworks—and be rooted in how those systems behave, not just how they are categorized or intended to function.

### v2

@@ Abstract
Lay Abstract:
We argue that recently developed military weapon and command systems using frontier AI are a categorically new type of weapon system with unique risks and regulatory needs. These systems have been deployed internationally in growing numbers and variety in recent years, and existing regulation and oversight mechanisms are insufficient and nonspecific. We discuss the importance of technical AI researchers contributing to this under-discussed and fast moving topic, and we outline specific definitions and policy proposals to aid future debate.
@@ 1 Introduction
Artificial intelligence (AI), also called machine learning (ML), is embedded in a growing number of deployed weapon systems. These include uncrewed aerial vehicles, uncrewed surface vessels and submarines, and battlefield coordination platforms that assist with targeting and decision support [demay2022alphadogfight, rivera2024escalation, caballero2024large, gusterson2024s, Posey2024, zie2025, rosenbach2025autonomous, horowitz2021speed, Lushenko2023-or, scharre2023four, schneider2024hand]. The rapid deployment of these AI-augmented systems has outpaced existing regulatory frameworks. Governance models for conventional lethal autonomous weapon systems (LAWS)—typically rule-based and non-adaptive—were not designed to address challenges posed by modern AI [horowitz2021speed, Lushenko2023-or, scharre2023four, schneider2024hand]. These challenges include opacity (hard-to-interpret decision processes), adaptivity (behavioral shifts in response to new inputs), and post-deployment drift (performance degradation over time) [rivera2024escalation, simmons2024ai, Lamparth2024-qk, atalan2025critical].
Yet in both policy and AI communities, distinctions between conventional LAWS and modern AI-powered systems remain poorly defined—and are absent from existing treaties, principles, and deployment norms. These systems may misclassify targets under unfamiliar conditions, escalate conflicts due to unpredictable outputs, or obscure human accountability through black-box decisions—all while using AI models originally developed in civilian contexts (see Table 1).
As these systems become more central to warfare, civilian AI advances gain national security relevance, and governments may restrict the movement of people, research, and AI services to gain military advantage [simmons2024ai, shah2024unit]. Given the risks that reckless development of AI weapons pose, we take the position that modern AI-powered autonomous weapon systems—what we call AI-LAWS—require distinct and novel regulations to avoid harms to AI research and the AI industry. Effective regulation must consider the underlying ML technologies, and AI experts from diverse fields must be directly involved in the regulatory process.
To motivate this, we first review the current state of military AI development and policy in section 2, highlighting the range of systems publicly announced and already deployed. Next, we introduce a functional definition of AI-LAWS based on AI involvement in targeting, advising, or decision-making (section 3), which we use to identify systems that pose novel risks and require distinct regulatory consideration. Lastly, we propose several technically-informed policy directions grounded in this definition to address the risks AI-LAWS pose to both AI research and global security (section 4).
@@ 2 Background
We first survey deployed and in-development systems that meet our criteria for AI-LAWS (section 3), with examples listed in Table 2. We then examine the unique risks these systems pose in subsection 2.2, and explain why AI expert involvement is essential in their development, use, and regulation. These risks are not theoretical—they are structural features of how AI-LAWS are developed and behave, already visible in deployed systems [king2025ai]. Performance may degrade in new environments, behaviors may escalate tensions, and lethal decisions may be made by models no one fully understands. We summarize these risks in Table 1. Finally, we review existing regulatory efforts and alternate positions in subsection 2.3, and highlight why they fail to address risks specific to AI-LAWS.
@@ 2.1 Existing AI-LAWS
AI-LAWS are already deployed worldwide. Systems using AI to make real-time targeting or coordination decisions have already been demoed and deployed across air, land, sea, and command domains. Table 2 highlights examples with documented AI capabilities, from drone “swarms” to battlefield planning platforms. Many of these systems rely on models and tools developed in civilian research settings. Yet, researchers often remain unaware that their work may shape combat outcomes—or be deployed without rigorous validation or constraint [shah2024unit, Lushe2024]. Some AI-LAWS have already seen combat [shah2024unit, scharre2023four]. The Russian Lancet loitering munition, for example—a drone that independently identifies and crashes into targets—reportedly includes AI-based autonomous targeting capabilities and has seen wide deployment in Ukraine [faragasso2023, armyrecognition2025, rassler2025horizon]. AI-LAWS span a wide range of platforms, from disposable drones to naval vessels and advanced aircraft [Adam2024] (Table 2). These systems are being developed by both major and minor powers—with many systems being actively deployed [Wareham2020, Bode2023-yf, shah2024unit]. For example, the Estonian Milrem THeMIS UGV, which includes autonomous navigation and potentially targeting, has seen battlefield use in Ukraine [forces2025, armytech2025]. Given the diversity of actors, platforms, and missions, even well-intentioned policies will fall short unless they account for the varied ways AI models are used across military operations [scharre2023four, grand2023artificial].
Table 1: Example risks introduced by AI-LAWS that meet our oversight criteria. Oversight must be grounded in technical behavior—not system intent—and informed by expertise from the AI research community. This list of risks is not intended to be comprehensive.
@@
Undetected failures in real-world use Gaps in validation and overtrust in AI outputs make it harder to detect failures during deployment—e.g. in stressful or unfamiliar contexts. Forest-trained targeting AI model misidentifies vehicles in deserts [garcia2022drivers, barassi2024toward], and commanders defer to faulty AI planning advice [kreps2023exploring, holbrook2024overtrust].
Opacity and black-box decision-making AI-based targeting systems are difficult for human operators to understand or audit, causing unpredictable behavior in combat. AI-enabled engagement system selects lethal target based on faulty sensors the operator cannot override in time, leading to unintended casualties [Lamparth2024-qk].
Erosion of AI research freedom and scientific exchange Classified funding streams and dual-use constraints limit openness, autonomy, and international collaboration. University AI lab begins operating under military classification rules, limiting publication and/or international collaboration [Daniels2019-ib, simmons2024ai, kania2021myths].
Channeling AI expertise towards military applications Civilian AI researchers are recruited into military programs, sometimes without clear disclosure or opt-out mechanisms. University scientists find their programs absorbed into a defense-funded lab, shifting away from open-source AI research [ASPI_2025, Daniels2019-ib].
Acceleration of arms races and regional or global instability Widespread access to AI-LAWS lowers barriers to conflict escalation. Deploying drone swarms without adequate testing/safeguards, risking conflict escalation [Geist2016-kn, ding2024technology].
Table 2: Examples of real-world AI-enabled military systems with publicly documented unmanned or frontier AI-centered capabilities. Includes platforms used for targeting, coordination, or decision support across air, land, sea, and command domains. Status reflects 2019–2025 public data: In Dev = In development, Demo = Demonstrated, Deployed = In deployment.
@@
18th Airborne Corps System [kriegler2020artificial] U.S. Army U.S. - Deployed
Lavender [gusterson2024s] IDF Israel - Deployed
ChatBIT [pomfret-2025] PLA China - Demo
@@ 2.2 AI-LAWS present unique risks
AI-LAWS create distinct risks across military, geopolitical, and institutional domains. AI-LAWS pose novel risks across three overlapping domains—military [scharre2023four], international relations [schneider2024hand, rivera2024escalation, Bode2023-yu], and institutional (e.g., universities, tech companies) [simmons2024ai, shah2024unit, Kreps2025-vm, ryseff2022exploring]—that differ from those associated with conventional LAWS. First, poorly validated or overtrusted systems can cause battlefield failures—including misidentification, friendly fire, brittle AI-human coordination, and performance gaps between exercises and combat [holbrook2024overtrust, sambasivan2021everyone, Kyiv2024-tq, fodorbenchmark, Biddle2024-ai]. Second, unpredictable or escalatory behavior can increase the risk of interstate conflict, especially when AI systems are deployed without adequate testing or validation [palantir2023, Lamparth2024-qk, rivera2024escalation, holbrook2024overtrust]. Third, the rapid pipeline from AI research to military use poses risks to scientific openness, research trajectories, and the stability of international norms [sarkis2023, scharre2023four, simmons2024ai, Lushe2024]. We summarize these risks in Table 1.
@@
In the military domain, the need to validate weapons performance during peacetime is not new—but the characteristics of modern AI make it more difficult than for conventional systems. Without additional validation, AI-LAWS are more likely to underperform in conflict, often in unintuitive ways. For example, “brittleness” is a well-documented issue in both academic and industrial AI, especially as systems become more advanced and multifunctional. It refers to the often rapid decline in performance when AI is exposed to contexts outside its training data—including variations a human would expect to be irrelevant [lohn2020estimating, druce2021brittle]. For example, an AI-LAWS trained on diverse biomes but validated only in temperate European environments may underperform in tropical or East Asian settings. This risk heightens when iterative tuning focuses on a narrow deployment context, causing unintentional overfitting to the validation setting. As a result, performance suffers in environments the system was trained on but never field-tested in. Such failures can also be difficult to detect— the AI models AI-LAWS rely on are partially opaque, and their bases of decisions may not be interpretable or auditable with high confidence [horowitz2021speed, caballero2024large, atalan2025critical].
@@
For the second domain, beyond battlefield performance, the development and spread of AI-LAWS may destabilize international relations in subtle and unpredictable ways. For example, the unpredictability of modern AI systems can create novel conflict risks if an AI-LAWS is deployed in a diplomatically tense region [Lamparth2024-qk, rivera2024escalation]. An AI-LAWS patrolling a border may protect soldiers, but could misinterpret warning shots as hostile and initiate or recommend return fire, exacerbating tensions that could escalate to conflict. Even without deployment, uncertainty over AI-LAWS battlefield impact has fueled arms-race dynamics, prompting major investments by multiple nations [Bode2023-yu, Adam2024]. These investments risk becoming a security dilemma that encourages a competitive arms race. The problem is compounded by technical uncertainty around AI-LAWS capabilities and timelines, and the need to test systems in real world combat [Kyiv2024-tq, Flynn2020-rk, sambasivan2021everyone]. Lacking better insight, policymakers may overestimate their impact to avoid under-investing, while also seeking conflict to test them in, as is the case in Ukraine. These dynamics raise the risk of escalation even before AI-LAWS are deployed.
@@
Lastly, AI-LAWS pose risks to academic and civilian AI freedoms and norms. Modern AI is distinctive in that the gap between civilian and military applications is unusually small. As a result, civilian advances are easily repurposed for military use, which has national and global security implications. This incentivizes governments to overregulate the process and dissemination of AI research and models, e.g., through publication blackouts and travel restrictions on AI researchers [simmons2024ai]. As civil-military barriers erode, regulatory oversight, export restrictions, and international controls may further limit civilian AI development. Some restrictions—such as controls on AI hardware exports—are already in place, and similar constraints on software and algorithms are likely to follow [BIS2024aiweights, Reuters2020chinacontrols, EUAIACT2024]. These risks are amplified by uncertainty. When policymakers lack clarity about AI’s national security implications, they may over-regulate to hedge against worst-case scenarios. These governance failures are addressed in the next section.
@@ 2.3 Existing policy efforts do not address unique AI-LAWS risks
Despite broad policy attention, current governance frameworks do not account for the technical risks posed by modern AI systems. Nations continue to field military AI, yet no binding standards exist to govern how the systems are designed, validated, or deployed [palantir2023]. Most existing LAWS and AI safety frameworks emphasize broad principles—such as human oversight or bias minimization—but lack concrete mechanisms and metrics tailored to how these systems function in practice.
Alongside these awareness issues, the AI research community has been largely isolated from military governance conversations. Initial regulatory efforts were led by defense officials, with limited technical input from the scientific community [pramudia2022china]. Today, oversight of military AI is debated at institutional, national, and international levels [HD2021, grand2023artificial]—yet those best equipped to evaluate system behavior remain underrepresented. Militaries worldwide have expressed concern over lacking AI expertise to update doctrine and regulation as technology evolves [shah2024unit, Anthony_Pfaff2023-sr, ukresponse2024, soare2023european]. This disconnect creates both risks and opportunities. To meaningfully address AI-specific failure modes, researchers must be at the table [cha2024towards]. AI researchers who seek to shape policy must engage with military actors beyond their usual academic spheres [Lushe2024, shah2024unit, WinnUnknown-ji]. Engagement is essential to ensure that governance reflects real-world system behavior—not just design intentions or policy abstractions.
Regulatory frameworks have not kept pace with failure modes such as opacity, drift, and escalatory misuse [Adam2024]. Technical and policy communities alike struggle to establish thresholds for oversight, especially as systems grow more adaptive, autonomous, and complex. Current certification procedures are rooted in conventional systems and offer little clarity on AI-specific metrics or standards [palantir2023, DoD2024, ukresponse2024]. These gaps are compounded by geopolitical concerns: policymakers may resist unilateral regulation for fear of losing strategic advantage in advancing AI technologies [Bode2023-yu], and international progress has been limited by misaligned priorities and inconsistent institutional mandates [chauhan2024artificial].
Efforts to define AI-LAWS—in the United Nations [Bode2023-yu], Track II forums [kraft2022dissent, Kahl2024-fc], or REAIM [jon2024scientific]—have struggled to yield useful standards [horowitz2016ban]. Technical experts remain underrepresented in these discussions, and proposed definitions often hinge on extreme thresholds like “full autonomy” or “superhuman learning” [pramudia2022china]. One influential definition, for example, requires that a system expand its capabilities beyond human expectations without any human intervention [Bode2023-yu]. While well-intentioned, such thresholds exclude most deployed and near-term systems—and thereby prevent oversight of the systems already in use.
@@ 3 Oversight Criteria for AI-LAWS, Grounded in System Behavior
Effective oversight of AI-LAWS requires moving beyond labels like “autonomous” or “general intelligence” and toward criteria grounded in how systems actually function. Targeting models can misidentify civilian vehicles in unfamiliar terrain [garcia2022drivers, Lamparth2024-qk]. Autonomous patrols may escalate based on ambiguous signals, without human authorization [rivera2024escalation, Lamparth2024-qk]. Command-advising models may drift over time in ways operators cannot detect [caballero2024large, Biddle2024-ai]. These risks are not hypothetical—they mirror failures already observed or anticipated in real-world systems (see Table 1). Oversight must begin with how systems behave, not how they are labeled.
Current frameworks often hinge on vague descriptors—like “human-in-the-loop” or “fully autonomous”—that fail to reflect how systems behave in practice. For example, a loitering munition may shift between human-authorized and AI-executed modes based on mission context [Bode2023-yf]. Command-advising systems may shape lethal decisions without directly executing them [Lushenko2023-or, caballero2024large]. Despite posing real governance challenges, such systems risk not qualifying as AI-LAWS under strict definitional thresholds—even when they should.
@@
  • Criterion 1: Use of AI/ML methods (e.g., neural networks) that are necessary for system function, and that pose AI-specific risks such as misidentifying targets, escalating unpredictably, drifting after deployment, or degrading in unfamiliar environments due to limited generalization beyond the training data distribution.
@@
One of the main barriers to regulating conventional LAWS has been disagreement over definitions [scharre2015autonomy, horowitz2016ban]. According to U.S. submissions to the UN, LAWS—sometimes called “killer robots”—are defined as systems that “can identify, select, and engage targets with lethal force without further intervention by an operator” [Unoda2023-bu]. Critics note that this definition is overly broad, potentially encompassing naval mines and heat-seeking missiles. The radar-seeking Israeli Harpy loitering munition, deployed in the late 1980s, exemplifies early LAWS already covered by prior regulations as historical models behave predictably with simple tracking functions [scharre2015autonomy]. Recent advances in AI research have significantly increased the combat power and autonomy of AI-enhanced systems [makridakis2017forthcoming, Scharre2018-oq, schneider2024hand, caballero2024large, Posey2024, shah2024unit]. As AI capabilities were added to LAWS, it became increasingly necessary to distinguish conventional automated weapons from AI-LAWS [Bode2023-yf, grand2023artificial, DoD2024]. While AI has long been used in simple forms for military tasks [grewal2010applications, Geist2016-kn], modern AI-LAWS employ far more advanced ML models, can operate along a continuum of modes from remote-controlled to fully autonomous, and present unique AI-specific risks and failure modes not seen in conventional LAWS.
@@
Oversight of military AI systems has traditionally centered around three narratives. While each raises valid concerns, none is technically grounded enough to address how deployed systems behave—particularly regarding opacity, drift, and brittleness when the domain of their use shifts. The first, rooted in humanitarian advocacy, calls for bans or strict human-in-the-loop constraints based on the principle of Meaningful Human Control (MHC) [Article2014-vf, scharre2015meaningful, Wareham2020]. The second emphasizes long-term existential risks from future artificial general intelligence (AGI) [bengio2025superintelligent, ienca2023don]. The third—most common in defense policy—treats AI-LAWS as a natural extension of conventional automation and maintains that existing frameworks, such as the U.S. DoD’s Directive 3000.09, are sufficient [DoD2024]. This approach, termed Appropriate Human Judgment (AHJ), is embedded in U.S. military doctrine [DoD2024, Ashley2024-wu]. It emphasizes procedural safeguards—such as system testing, operator training, and robust design—rather than requiring direct human control over every lethal force application. AHJ is often viewed as more pragmatic for increasingly complex systems—such as when operations require faster response times than humans can provide, involve cooperative autonomous swarms, or scale beyond what human oversight can track reasonably in real time [rosenbach2025autonomous].
@@
Yet all three narratives remain underdefined and difficult to operationalize. MHC lacks measurable criteria for what constitutes “meaningful” control in fast-moving or distributed contexts. AHJ, though more scalable, often fails to define metrics for what constitutes sufficient “judgment”—or to codify how to assess the safety of opaque, adaptive, or continually learning systems [DoD2024]. Critics note that AHJ can obscure accountability behind vague assurances of system design—especially when those assurances come from third-party contractors or defense tech firms [Lamparth2024-qk, rivera2024escalation, Lushe2024, Biddle2024-ai, Bode2021].
@@
Policy limitations have real consequences. Post-deployment drift, overreliance on black-box systems, brittle generalization, and institutional pressure to avoid challenging flawed AI outputs can all occur under either oversight model [Adam2024]. Command-advising systems may formally keep humans “in the loop,” yet users often defer to flawed recommendations—especially under time pressure or when authority to intervene is unclear [Hunter2022-bw, holbrook2024overtrust]. To be effective, any oversight framework—whether based on MHC, AHJ, or another model—must be paired with rigorous behavioral validation, grounded in AI research expertise. This includes task-specific performance benchmarks, failure audits, stress testing under domain shift, and post-deployment monitoring. These mechanisms do not replace high-level policy principles; they make them testable and enforceable. The U.K.’s 2024 guidance for military AI use offers a promising start, with proposals for AI-specific audits, operator certification, and red-line prohibitions [ukresponse2024]. However, as AI capabilities advance, fixed rules alone cannot keep pace. Governance must evolve alongside systems with flexible frameworks—and be rooted in how those systems behave, not just how they are categorized or intended to function.